OpenStudy (anonymous):

A line passes through the point (−2, 7) and has a slope of 8. Write an equation for this line.

OpenStudy (anonymous):

Do u want just the answer or do u want to find out how to get this answer?

OpenStudy (anonymous):

both please?

OpenStudy (anonymous):

Ok are you regular school or online?

OpenStudy (anonymous):

online

OpenStudy (anonymous):

You want to find the equation for a line that passes through the point (-2,7) and has a slope of 8. First of all, remember what the equation of a line is: y = mx+b Where: m is the slope, and b is the y-intercept To start, you know what m is; it's just the slope, which you said was 8. So you can right away fill in the equation for a line somewhat to read: y=8x+b. Now, what about b, the y-intercept? To find b, think about what your (x,y) point means: (-2,7). When x of the line is -2, y of the line must be 7. Because you said the line passes through this point, right? Now, look at our line's equation so far: . b is what we want, the 8 is already set and x and y are just two "free variables" sitting there. We can plug anything we want in for x and y here, but we want the equation for the line that specfically passes through the the point (-2,7). So, why not plug in for x the number -2 and for y the number 7? This will allow us to solve for b for the particular line that passes through the point you gave!. (-2,7). y=mx+b or 7=8 × -2+b, or solving for b: b=7-(8)(-2). b=23. The equation of the line that passes through the point (-2,7) with a slope of 8 is y=8x+23

OpenStudy (anonymous):

(-2,7). y=mx+b or 7=8 x -2+b, or solving for b: b=7-(8)(-2). b=23. sorry it showed "?????" before

OpenStudy (anonymous):

answer is y=8x+23 and i've explained the method before

OpenStudy (anonymous):

Thank you so much! You saved me big time :p

OpenStudy (anonymous):

np :)